) Which ion does NOT have a noble gas configuration in its ground state?

 

           A) Si4+

           B) Al3+

           C) Sn2+

           D) Se2-

           E) All of the above

 


1
Expert's answer
2021-05-27T07:34:08-0400

In Noble gas configuration must have 8 valence electrons either by gaining or losing electrons.

So selenium in not a noble gas in its ground state. answer D is correct.


Si4+ → +2 charge → loss of 4 electron → +2 +2 -4 = 0 or 8 valence electron → noble gas configuration

Al3+ → +3 charge  → loss of 3 electrons → 3 - 3 = 0 or 8 valence electrons → noble gas configuration

Sn2+ → +2 charge → loss of 2 electron →2-2 =0 or 8 valence electron → noble gas configuration

Se2- → +4 charge → add of 2 electron → 4+2 =6 so → not a noble gas configuration